OpenStudy (anonymous):

A 1.0mm diameter, 20cm long copper wire carries a 3.0A current. What is the potential difference between the ends of the wire?

OpenStudy (ybarrap):

$$ R=\rho\cfrac{l}{A} $$ source - http://en.wikipedia.org/wiki/Electrical_resistivity#Resistors_or_conductors_with_uniform_cross-section Where $$ A=\pi r^2~m^2\\ r=\cfrac{.001}{2}~m\\ l=0.20~m\\ \rho=16.78\times 10^{-9}~\Omega\cdot m $$ source - http://en.wikipedia.org/wiki/Electrical_resistivity Potential Difference: $$ Ans. V=IR\\ \text{where }I=3.0~A $$ Make sense?